# Hoaxpedia: A Unified Wikipedia Hoax Articles Dataset

### v2

@@ 4 Experiments
We cast the problem of identifying hoax vs. legitimate articles as a binary classification problem, in whcih we evaluate a suite of LMs, specifically: BERT-family of models (BERT-base and large Devlin et al. (2019), RoBERTa-base and large Liu et al. (2019), Albert-base and large Lan et al. (2019)), as well as T5 (Base and Large) Raffel et al. (2020) and Longformer (Base) Beltagy et al. (2020). We use the same training configuration for the BERT-family of models, T5 models and Longformer, and set the generation objective as Binary classification for the T5 models. In terms of data size, we consider the three different scenarios outlined in Section 2 (2x, 10x and 100x negative examples). This approach naturally increases the challenge for the classifiers. The details about the data used in different settings are given in Appendix A.
In addition to the three different settings for positive vs negative ratios, we also explore how much text is actually needed to catch a hoax, or in other words, are definition sentences in hoax articles giving something away? This is explored by running our experiments on the full Wikipedia articles, on one hand, and on the definition (first sentence alone), on the other. This latter setting is interesting from a lexicographic perspective because it helps us understand if the Wikipedia definitions show any pattern that a model could exploit. Moreover, from the practical point of view of building a classifier that could dynamically “patrol” Wikipedia and flag content automatically, a definition-only model would be more interpretable (with reduced amiguity and focusing on core meaning/properties of the entity) and could have less parameters (handling smaller vocabularies, and compressed knowledge), which would have practical retraining/deployment implications in cost and turnaround.
@@ 5.1 Ablation: Effect of Definitions on Classifying Hoaxes
We run a data ablation test on the full-text split of the dataset to find the impact of definition sentences. To this end, we remove the first sentence of each article, and replicate the “full text” classification experiment, focusing on RoBERTa-Large, the most consistent model. The results of this ablation experiment are shown in Table 2,suggesting that F1 goes down about 2% for the positive class when the definition sentence is missing. This shows that definitions show critical information about entities and events in Wikipedia, but often are not the place where hoax features would emerge, and therefore removing them from the full text doesn’t change much the story.
@@
RoBERTaL Ft 1:2 0.83 0.80 0.82
RoBERTaL Ft 1:10 0.82 0.71 0.76
RoBERTaL Ft 1:100 0.67 0.51 0.58
Table 2: Performance of RoBERTa-Large on Binary Classifications Without Definition Sentences in Articles (with Fulltext Hoax:Real ratio in Settings column)
@@ 6 Conclusion and Future Work
We have introduced HoaxPedia, a dataset containing hoax articles extracted from Wikipedia, from a number of sources, from official lists of hoaxes, existing datasets and Web Archive. We paired these hoax articles with similar real articles, and after analyzing their main properties (concluding they are written with very similar style and content), we report the results of a number of binary classification experiments, where we explore the impact of (1) positive to negative ratio; and (2) going from the whole article to only the definition. This is different from previous works in that we have exclusively looked at the content of these hoax articles, rather than metadata such as traffic or longevity. For the future, we would like to further refine what are the criteria used by Wikipedia editors to detect hoax articles, and turn those insights into a ML model, and explore other types of non-obvious online vandalism.
@@ 7 Ethics Statement
Appendix A Appendix: Dataset Details
@@
1Hoax2Real Definition Train 426 206 632
@@
1Hoax2Real Full Text Train 456 232 688
@@
1Hoax10Real Definition Train 2225 203 2428
Test 940 104 1044
1Hoax10Real Full Text Train 2306 218 2524
Test 973 110 1083
1Hoax100Real Definition Train 20419 217 20636
Test 8761 82 8843
1Hoax100Real Full Text Train 22274 222 22496
Test 9534 106 9640
@@
Appendix B Appendix: Training Details
We train our models with the configuration given below. We use one NVIDIA RTX4090, one NVIDIA V100 and one NVIDIA A100 GPU to train the models.

### v3

@@ 5 Experiments
We cast the problem of identifying hoax vs. legitimate articles as a binary classification problem, in whcih we evaluate a suite of LMs, specifically: BERT family of models (BERT base and large Devlin et al. (2019), RoBERTa-base and large Liu et al. (2019), Albert-base and large Lan et al. (2019)), as well as T5 (Base and Large) Raffel et al. (2020) and Longformer (Base) Beltagy et al. (2020). We use the same training configuration for the BERT-family of models, T5 models and Longformer, and set the generation objective as Binary classification for the T5 models. In terms of data size, we consider the three different scenarios outlined in Section 3 (2x, 10x and 100x negative examples). This approach naturally increases the challenge for the classifiers. The details about the data used in different settings are given in Appendix A.
In addition to the three different settings for positive vs negative ratios, we also explore how much text is actually needed to catch a hoax, or in other words, are definition sentences in hoax articles giving something away? This is explored by running our experiments on the full Wikipedia articles, on one hand, and on the definition (first sentence alone), on the other. This latter setting is interesting from a lexicographic perspective because it helps us understand if the Wikipedia definitions show any pattern that a model could exploit. Moreover, from the practical point of view of building a classifier that could dynamically “patrol” Wikipedia and flag content automatically, a definition-only model would be more interpretable (with reduced ambiguity and focusing on core meaning/properties of the entity) and could have less parameters (handling smaller vocabularies, and compressed knowledge), which would have practical retraining/deployment implications in cost and turnaround.
@@ 6.1 Ablation: Effect of Definitions on Classifying Hoaxes
We run a data ablation test on the full-text split of the dataset to find the impact of definition sentences. To this end, we remove the first sentence of each article, and run the fulltext classification experiment, focusing on RoBERTa-Large, the most consistent model. The results of this ablation experiment are shown in Table 4, suggesting that F1 goes down about 2% for the positive class when the definition sentence is missing. This shows that definitions show critical information about entities and events in Wikipedia, but often are not the place where hoax features would emerge, and therefore removing them from the full text doesn’t change much the story.
@@
RoBERTaL 1H2R 0.83 0.80 0.82
RoBERTaL 1H10R 0.82 0.71 0.76
RoBERTaL 1H100R 0.67 0.51 0.58
Table 4: Performance of RoBERTa-Large on Binary Classifications Without Definition Sentences in Articles (with Fulltext hoax to real ratio in Settings column)
7 Large Language Models and HoaxPedia
We explore the capabilities of open source Large Language Models (LLM) in detecting hoax articles through our proposed dataset. We select Llama2-7B and 13B Touvron et al. (2023), Llama3-8B Dubey et al. (2024), and Mistral-7B models Jiang et al. (2023) for the experiments, and the prompts used are given in the Appendix C. We take three experimental settings: perplexity, instruction-tuning, and supervised fine-tuning. The settings and their results are described below.
Perplexity Experiments: We consider perplexity as an indicator for LLMs to predict distribution of of Hoax and legitimate articles, with the hypothesis that LLMs will have difficulty for predicting contents of hoax articles, resulting in higher perplexity. We test the LLMs in both definition and fulltext settings. The average perplexity results for both settings are shown in Figure 6, revealing that there is a significant difference between perplexity of hoax and legitimate articles in both settings. This suggests that LLMs struggle to predict distribution of Hoax articles.
(a) Average perplexity scores for fulltext
(b) Average perplexity scores for definition
Figure 6: Average Model based perplexity scores for Fulltext and Definition Settings
Prompting: For prompting, we consider zero-shot and few-shot prompt settings. The prompts are given in Appendix C, and the input setting are for both definition and fulltext. We report the results for Macro-F1 scores. The results are reported in Table 5, showing that the Mistral model show consistency in performance across the splits, albeit it is not the best model. Llama2-13B models perform the best for both the settings (definition and fulltext). The results also show that the performance difference between the definition and fulltext setting is marginal, as opposed to fine-tuned Language Models in Table 3.
Model Name Zero-shot Few-shot
Definition Fulltext Definition Fulltext
Llama2-7B 0.48 0.50 0.51 0.52
Llama2-13B 0.57 0.58 0.59 0.59
Llama3-8B 0.33 0.40 0.35 0.40
Mistral-7B 0.53 0.56 0.54 0.58
Table 5: F1 score for Prompting experiment in Zero and Few shot setting for Definition and Fulltext splits
Fine-tuning: We also fine-tune the models tested with perplexity and instruction tuning with HoaxPedia with supervised fine-tuning Touvron et al. (2023). The results of fine-tuning as Macro-F1 scores for both definition and fulltext setting are shown in Table 6. The results show significant improvement across all the settings for all the models. Llama3 shows most consistency across the settings, and also is the best model across the scenarios, with a performance improvement of more than 25% across the scenarios. However, the performance of the models in 1:100 hoax to legitimate scenario for all the models are around 50% F1 in both settings. Further analysis reveals that the models only predict the majority label, which can further attributed to imbalance in the dataset Powers (2020). A suitable weight adjusted training paradigm is required to handle the imbalance, but since it is not the scope of this work, we keep it for future work.
Model Definition Fulltext
1H2R 1H10R 1H100R 1H2R 1H10R 1H100R
Llama2-7B 0.76 0.47 0.49 0.66 0.48 0.47
Llama2-13B 0.80 0.48 0.50 0.60 0.63 0.50
Llama3-8B 0.80 0.48 0.50 0.83 0.67 0.50
Mistral-7B 0.71 0.55 0.49 0.68 0.53 0.49
Table 6: Macro-F1 score for LLM Fine-tuning in degrees of data imbalance for Definition and Fulltext Settings
@@ 8 Conclusion and Future Work
We have introduced HoaxPedia, a dataset containing hoax articles extracted from Wikipedia, from a number of sources, from official lists of hoaxes, existing datasets, and the Web Archive. We paired these hoax articles with similar legitimate articles, and after analyzing their main properties (concluding they are written with very similar style and content), we report the results of a number of binary classification experiments, where we explore the impact of (1) positive to negative ratio; and (2) going from the whole article to only the definition. This is different from previous work in that we have exclusively looked at the content of these hoax articles, rather than metadata such as traffic or longevity. For the future, we would like to further refine what are the criteria used by Wikipedia editors to detect hoax articles, and turn those insights into a ML model, and explore other types of non-obvious online vandalism.
9 Limitations
We present a new dataset named Hoaxpedia and associated baselines from a wide variety of language models / large language models in this work. Our study shows that these types of dataset can be helpful in the area of free text disinformation detection. However, there are some limitations to our work that we aim to address here. The sets proposed here are small, with only 311 positive examples (hoaxes), which can be attributed to the fact that we only collect the examples that are explicitly labeled as hoaxes, rather than articles under discussion for hoaxes. Additionally, in our experiments, we do not conduct further investigation for model behaviors such as performance improvement of Longformer models in the hardest setting (cf. Section 6, implementing a weight-adjusted training paradigm for fine-tuning of LLMs (cf. Section 7) etc. We leave these analysis in future work, as the scope of this work is to introduce this dataset and establish the baseline results. Finally, we do not compare the results with existing work, mainly with Kumar et al. (2016), since the approaches mentioned in existing work are metadata dependent, and our approach is based on article text, we argue that the results may not be comparable.
@@ 10 Ethics Statement
Appendix A Dataset Details
@@
1Hoax2legitimate Definition Train 426 206 632
@@
1Hoax2legitimate Full Text Train 456 232 688
@@
1Hoax10legitimate Definition Train 2,225 203 2,428
Test 940 104 1,044
1Hoax10legitimate Full Text Train 2,306 218 2,524
Test 973 110 1,083
1Hoax100legitimate Definition Train 20,419 217 20,636
Test 8,761 82 8,843
1Hoax100legitimate Full Text Train 22,274 222 22,496
Test 9,534 106 9,640
@@
Appendix B Language Model Training Details
We train our Language Models with the configuration given below. We use one NVIDIA RTX4090 to train the LMs, one NVIDIA V100 and one NVIDIA A100 GPU to train the LLMs.
@@ Appendix B Language Model Training Details
Appendix C Prompt for LLM in-context learning
The instruction prompt used for LLMs in their in-context learning with examples for few shot experiment are given below.
You are a helpful knowledge management expert and excel at identifying whether an input Wikipedia article is a hoax or not. Wikipedia defines a hoax as ‘a deliberately fabricated falsehood made to masquerade as truth’. You take an Wikipedia article as input and return with the label citing hoax(Label 1) or real(Label 0) based only on the text of the article. Given an article from Wikipedia, your task is to analyze the article text to identify if the article is hoax or real. The Hoax and real articles are defined as follows: Hoax: An article that is deliberately fabricated falsehood made to masquerade as truth. Real: An article which contains information about an existing entity and are not fabricated. Your output should be a JSON dictionary with label that you found. Here are the possible labels with what they mean: 0 : The article is real article. 1 : The article is a hoax article. Your input will be in the following format: INPUT: { Text: <Article text> } OUTPUT: { Label: <One of the label from the possible labels: 0 and 1, where 0 is real article and 1 is hoax article.> } Please respond with only the JSON dictionary containing label. You are instructed strictly to return output only in the format given above, nothing else. No yapping.
Here are the examples used in few-shot experiments.
Example 1: INPUT: { Text: Albion Dauti (born May 31, 1995 in Caracas, Venezuela) is a Venezuelan telenovela actor and presenter. Albion Dauti was born in Caracas, Venezuela. He studied acting at the Faculty of Arts in Caracas. In 2010 he started working as a presenter in Venevision. } OUTPUT: { Label: 1 } Example 2: INPUT: { Text: Michelle Madhok (born May 26, 1971) is the Founder and CEO of White Cat Media Inc. - DBA SheFinds Media, parent company of online shopping publication SheFinds. com and MomFinds. com. She writes a weekly style column for New York’s Metro newspaper and appears regularly on Fox News Channel, The Today Show, and The Tyra Banks Show. } OUTPUT: { Label: 0 }